OpenStudy (anonymous):

What is the geometric mean of 175 and 7?

OpenStudy (anonymous):

Multiply and take squareroot. The Geometric Mean, \(GM\) is defined: \[GM=\sqrt[n]{x_1x_2\cdots x_n}\]

Parth (parthkohli):

Geometric mean of a and b is \(\sqrt{ab}\)

Parth (parthkohli):

Geometric mean of a,b,c is \(\sqrt[3]{abc}\)

Parth (parthkohli):

etc. etc.

OpenStudy (anonymous):

Yup. It's precisely as Parth and I said. I just wrote it using subscripts.

OpenStudy (anonymous):

Well, the answer choices are 55, 35, 45, and 40.

OpenStudy (anonymous):

Are you allowed a calculator? If so, use it.

OpenStudy (anonymous):

No calculator allowed. =\

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sqrt{175 \times 7} }\) Find the value :)

OpenStudy (anonymous):

Ok, then you have to find the prime factorization of \(175\).

OpenStudy (amistre64):

square the answers til you get the underradical ....

OpenStudy (anonymous):

Amistre's method also works.

Parth (parthkohli):

Yeah, first multiplying them and then finding the square root wouldn't make it easier.

OpenStudy (amistre64):

a sqrt by hand is about as hard is long division; but if given answers to compare to; its easier to square the answers :)

OpenStudy (anonymous):

In most cases, yes. I think this is not as hard, though.

OpenStudy (anonymous):

Okay, answer is 35, thankyou.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sqrt{175 \times 7} = \sqrt{5 \times 7 \times 5 \times 7 } }\)

Parth (parthkohli):

Yes, it is 35.

OpenStudy (anonymous):

\(175\) has a relatively simple factorization. (I'm referring to algorithmic complexity.)

OpenStudy (amistre64):

we knew you could do it ;)

OpenStudy (anonymous):

:D

Parth (parthkohli):

lol

Parth (parthkohli):

The ones ending with 5 are easy because you can start by dividing 5 :)

OpenStudy (anonymous):

Okay, okay, i suck at math, pardon me.

Parth (parthkohli):

No you don't, because you got the answer :)

OpenStudy (anonymous):

Oh, no. You don't suck at math.

Parth (parthkohli):

Most of the people aren't able to get the answer, but you are, and easily :D

OpenStudy (amistre64):

this site is here to help you get better at math; and not just hand out free answers ... so you did remarkable well for sucking

Parth (parthkohli):

Our second best member says so ^ you must feel proud :D

OpenStudy (anonymous):

LOL, Amistre. Nicely put.

OpenStudy (anonymous):

I can write a flawless MLA format paper on phisics and astronomy or anything else you'd want. Give me 10th grade math and my brain turns into mush.

OpenStudy (anonymous):

Everyone has a weak area. Just because it's a weak area doesn't mean they suck, though. A lot of mathematicians consider themselves weak in composition and grammar. But you don't see them thinking they suck--nor using composition or grammar which sucks. :P You just have to be more confident in yourself and keep trying your best. You are doing fine as far as I can tell.

Parth (parthkohli):

I suck at calculus :P

OpenStudy (anonymous):

Well, i am in novanet for math, so aperantly it is not my forte.

OpenStudy (anonymous):

Randi, feel free to ask any questions on OS. There is a diversity of expertise here. I wish you luck in your maths.

OpenStudy (anonymous):

Yes, sense'.